# A semi-parametric marginalized dynamic conditional correlation framework

### v3

@@ Abstract
We develop a novel multivariate semi-parametric framework for joint portfolio Value-at-Risk and Expected Shortfall forecasting. Unlike existing univariate semi-parametric approaches, the proposed framework explicitly models the dependence structure among portfolio asset returns through a marginalized dynamic conditional correlation (DCC) parameterization. To estimate the model, a two-step procedure based on the minimization of a strictly consistent scoring function derived from the Asymmetric Laplace distribution is developed. This procedure allows to simultaneously estimate the marginalized DCC parameters and the portfolio risk factors. The performance of the proposed model in risk forecasting and portfolio allocation is evaluated by means of a forecasting study on the components of the Dow Jones index for an out-of-sample period from December 2016 to September 2021. The empirical results support effectiveness of the proposed framework compared to a variety of existing approaches.
Keywords: multivariate; semi-parametric; Value-at-Risk; Expected Shortfall; forecasting; portfolio allocation.
@@ 1 Introduction
Value-at-Risk (VaR) and Expected Shortfall (ES) play a central role in the risk management systems of banks and other financial institutions. For more than two decades, VaR has been the official risk measure adopted worldwide by financial intermediaries operating in the global financial system. Nevertheless, VaR has some important theoretical limits. First, VaR cannot measure the expected loss for extreme (violating) returns. In addition, it can be shown that VaR is not always a coherent risk measure, due to failure to match the subadditivity property. For these reasons, the Basle Committee on Banking Supervision proposed in May 2012 to replace VaR with the ES (Artzner 1997; Artzner et al. 1999). The prominent role of ES in the new international risk measurement practice is clearly illustrated by the official standards for determining minimum capital requirements for market risk published on the BIS institutional website (Basel Committee on Banking Supervision, 2019, see e.g. section MAR33) and effective from January 2022. Thus, in recent years ES has been increasingly employed for tail risk measurement. Analytically, ES is defined as the expectation of the return conditional on the VaR being exceeded and, unlike VaR, it is a coherent measure and “measures the riskiness of a position by considering both the size and the likelihood of losses above a certain confidence level” (Basel Committee on Banking Supervision, 2013).
Our paper focuses on the daily forecasting of VaR and ES on the lower/left tail. In line with the Basel III Accord, our empirical analyses use the common $\alpha=2.5\%$ probability level, which is the standard reference value for ES forecasting.
@@
We contribute to the literature on portfolio risk forecasting by proposing a novel class of semi-parametric marginalized multivariate GARCH (MGARCH) models that jointly generate portfolio VaR and ES forecasts, using multivariate information on the returns of portfolio constituents as input. Our main reference model is given by a marginalized version of the dynamic conditional correlations (DCC) model (Engle, 2002) whose parameters are estimated by minimization of a negative AL log-likelihood as in Taylor (2019), without requiring the formulation of restrictive assumptions on the conditional distribution of portfolio returns.
In the remainder, for short we will refer to this model as the DCC-AL model. The proposed framework can be easily extended to other parameterizations from the multivariate GARCH literature such as the corrected DCC (cDCC) of Aielli (2013) or dynamic covariance models such as the BEKK model of Engle and Kroner (1995). Our preference for the DCC parameterization is motivated by the flexibility of this specification and its widespread diffusion among practitioners  [^2] We also developed our framework assuming a cDCC model for correlation dynamics, cDCC-AL for short. However, consistently with previous findings, the cDCC-AL and DCC-AL models turned out to give very similar estimates, motivating our decision to keep the mathematically simpler DCC-AL as the main reference model. Results for the cDCC-AL model are available upon request.. At the same time, the AL scoring function can be replaced in a straightforward manner by any other strictly consistent loss for the pair (VaR, ES) such as the FZ0 used by Patton et al. (2019)
@@
There are two different ways to look at our model. Namely, the DCC-AL can be viewed either as a semi-parametric DCC model or as an ES-CAViaR for portfolio risk forecasting. It should be noted, however, that compared to each of these approaches, the DCC-AL is characterized by some distinctive and innovative features.
@@
Namely, compared to the standard DCC, the structure of the DCC-AL is made up of three components where the first two, pertaining to the specification of volatility and correlation dynamics, are in common with the usual DCC model while the third one, pertaining to portfolio risk estimation, is specific to our model. We refer to our proposed model as a marginalized semi-parametric DCC model since the risk model, while still using multivariate information on asset correlations, is fitted to the time series of univariate portfolio returns $r_{t}(\mathbf{w})$, for a predefined choice of portfolio weights $\mathbf{w}$. Furthermore, in the DCC model the dynamic conditional correlation parameters are usually estimated by Gaussian Quasi Maximum Likelihood (DCC-QML) through the maximization of a multivariate normal quasi-likelihood function. Then the portfolio VaR and ES could be ex-post obtained through the filtered historical simulation (HS) approach by calculating the sample quantiles and tail averages of the standardized returns (Francq and Zakoïan, 2015). In contrast, the DCC-AL optimizes a univariate risk-targeted strictly consistent loss function for portfolio VaR and ES, e.g. the AL based joint loss function, allowing portfolio VaR and ES to be estimated directly. This procedure also allows the simultaneous estimation of dynamic correlation parameters and portfolio risk factors.
@@
Compared to the standard ES-CAViaR framework, the DCC-AL model extends it to a multivariate setting, where it allows to obtain joint estimates of portfolio VaR and ES that explicitly take into account the cross-sectional correlation structure of portfolio returns. From a risk forecasting perspective, the DCC-AL model can then be viewed as an ES-CAViaR that takes multivariate input information on the asset variance and covariance matrix, which is marginalized to obtain univariate VaR and ES measures for portfolio returns. However, unlike univariate semi-parametric approaches to risk forecasting, the underlying multivariate GARCH structure allows the DCC-AL model to be used for a wider range of applications, including portfolio allocation and hedging.
@@
The empirical performance of the DCC-AL model is assessed by means of an application to a panel of 28 assets included in the Dow Jones index. Our findings can be summarized as follows. When forecasting risk for an equally weighted portfolio, DCC-AL models are competitive with state-of-the-art univariate semi-parametric approaches and perform better than conventional DCC models. Next, we assess the performance of the proposed model in constructing global minimum ES portfolios. Under this respect, our results show that the DCC-AL model clearly outperforms a benchmark equally weighted allocation strategy. We also find evidence that in a turbulent period centred on the outbreak of the COVID-19 crisis in 2020, the DCC-AL model leads to less risky portfolios compared to existing DCC specifications.
The paper is structured as follows. Section 2, after providing a brief review of the existing semi-parametric univariate risk forecasting approaches, presents the proposed class of semi-parametric marginalized DCC models, while their estimation procedure is illustrated in Section 3. The finite sample properties of the estimators are investigated in Section A via an extensive Monte Carlo simulation study. Section 5 discusses the results of an empirical application to risk forecasting and portfolio optimization for the constituents of the Dow Jones index. Section 6 concludes the paper.
@@ 2.1 Description of the environment
In the remainder of Section 2 and 3, considering that we are focusing on risk forecasting at the $\alpha=2.5\%$ level for a pre-defined portfolio composition $(\mathbf{w})$, the following notational conventions will be adopted: $Q_{t,\alpha}(\mathbf{w})\equiv Q_{t}$, $\text{ES}_{t,\alpha}(\mathbf{w})\equiv\text{ES}_{t}$, $q_{\alpha}(\mathbf{w})\equiv q$, $c_{\alpha}(\mathbf{w})\equiv c$ and $r_{t}(\mathbf{w})\equiv r_{t}$.
@@ 2.2 Univariate semi-parametric approaches to portfolio risk forecasting
This allows a quasi-likelihood function to be built and maximised, given model expressions for $(Q_{t},\,\text{ES}_{t})$. Taylor (2019) supports the validity of this estimation procedure noting that the negative logarithm of the resulting likelihood function is strictly consistent for $(Q_{t},\,\text{ES}_{t})$ considered jointly, that is it fits into the class of jointly consistent scoring functions for VaR and ES developed by Fissler and Ziegel (2016). Along the same lines, Patton et al. (2019) investigate a class of semi-parametric models, including some observation driven models, whose parameters can be estimated minimizing a 0-degree homogeneous loss function (FZ0) still included in the same class. Gerlach and Wang (2020) extend the framework in Taylor (2019) by incorporating realized measures as exogenous variables, showing improved VaR and ES forecast accuracy.
As mentioned above, all these papers focus on univariate semi-parametric modeling approaches that, when applied to portfolio returns, do not explicitly assess the impact of cross-sectional correlations among assets. Although this issue has been extensively analyzed in the literature on parametric MGARCH models (Bauwens et al., 2006), to the extent of our knowledge, no attempts have been made to address it in a multivariate semi-parametric risk-targeted framework. In order to fill this gap, in the next section we propose a novel modeling strategy based on the use of what we call a semi-parametric marginalized DCC model.
2.3 A semi-parametric marginalized DCC model
The analytical expressions for portfolio VaR and ES provided in Equation (3) make evident that the VaR and ES dynamics are driven by those of the conditional variance and covariance matrix $H_{t}$. In this section we describe a semi-parametric framework for assessing the impact that conditional correlations among portfolio components and their individual volatilities have on portfolio risk forecasts. The dynamics of $H_{t}$ can be modelled using a wide range of specifications from the MGARCH literature. Without any loss of generality, due to its flexibility and widespread diffusion among practitioners, our proposed modeling approach builds on the DCC model of Engle (2002) and shares the same volatility and correlation dynamics. However, compared to the standard DCC, our framework incorporates an additional step related to the specification of portfolio VaR and ES.
We refer to the proposed model as a marginalized semi-parametric DCC-AL model. We call it marginalized since the risk model is fitted to the time series of univariate portfolio returns $r_{t}(\mathbf{w})$, for a given set of portfolio weights $\mathbf{w}$. The semi-parametric nature of the model derives from the fact that, as it will be later discussed in Section 3, its coefficients are fitted minimizing a jointly consistent loss for $Q_{t}$ and $\text{ES}_{t}$ without assuming the parametric distribution of returns. In particular, we use the negative of the AL log-likelihood in Equation (4) as the loss function.
@@ 2.3 A semi-parametric marginalized DCC model
Given the assumptions on $a$ and $b$, no identifiability issue arises and only four coefficients need to be estimated in the second stage of the DCC-AL model: $q$, $a$, $b$ and $\gamma_{0}$. The parameter estimate $q$ (VaR factor) allows us to directly estimate the quantile of the error distribution in the model for portfolio returns. Since we have defined $c^{2}/q^{2}=1+\text{exp}(\gamma_{0})$, after estimating parameters $q$ and $\gamma_{0}$, the value of the ES factor $c$ can be immediately recovered.
@@ 3 Estimation
The estimation of the DCC-AL model can be performed by means of a two step procedure in the spirit of Engle (2002).
@@
In step 1, we use the Matlab 2023a “fmincon” optimization routine to estimate the ES-CAViaR-IG models. In step 2, the Matlab “MultiStart” facility is further employed for “fmincon” to improve the robustness of the estimation results. We use 5 separate sets of starting points, generating 5 local solutions, then the optimum set among these is finally chosen as the parameter estimates  [^4] Details on “MultiStart” are provided in the Matlab documentation https://au.mathworks.com/help/gads/multistart.html.
@@ 4 Simulation study
A simulation study has been conducted in order to assess the statistical properties of the two-step AL loss based estimation procedure discussed in Section 3. Our analysis has two main objectives. First, the study evaluates the bias and efficiency of the estimators of the DCC-AL parameters ($a,b,q,\gamma_{0}$). Second, we assess the the risk forecasting performances of the estimated DCC-AL model in an equally weighted portfolio, via comparing the one-step-ahead $2.5\%$ level portfolio VaR and ES forecast accuracy to the “true” simulated values.
@@
The DCC-AL model is then fitted to the time series of equally weighted $(\mathbf{w}=1/n)$ portfolio returns generated from the chosen DGP, for three different sample sizes $T\in\{2000,3000,5000\}$. Overall, matching 3 different distributional assumptions (including both spherical and non-spherical distributions) and sample sizes, 9 simulation settings have been obtained and, under each of these, 250 return series have been generated.
Conditional on the chosen error distribution and on the values of the correlation parameters $a$ and $b$, the true values of the other two parameters $q$ and $\gamma_{0}$ in the DCC-AL model can be then calculated as follows.
@@
Therefore, taking the multivariate Normal case as an example, in the DCC-AL the true value of $q^{2}$ is $\Phi^{-1}(\alpha)^{2}=(-1.96)^{2}=3.8415$. The true value for $c^{2}$ is $\left(\frac{\phi(\Phi^{-1}(\alpha))}{\alpha}\right)^{2}=(-2.3378)^{2}$, where $\Phi$ is the standard Normal CDF and $\phi$ is standard Normal Probability Density Function (PDF). Further, we have $(1+\text{exp}(\gamma_{0}))=\frac{c^{2}}{q^{2}}=\frac{(-2.3378)^{2}}{(-1.9600)^{2}}=1.4227$, thus $\gamma_{0}=\text{log}(1.4227-1)=-0.8610$. These true values of $a$, $b$, $\gamma_{0}$ and $q^{2}$ for the multivariate Normal distribution are shown in the True rows in Table A1. When $\mathbf{z}_{t}\overset{iid}{\sim}t_{n}(\mathbf{0},I_{n};\nu=10)$, the true values of $q$, $c$ and $\gamma_{0}$ are obtained through a similar approach, replacing the analytical formulas for $q$ and $c$ with the equivalent expressions for an univariate $t$ distribution.
In the $nst$ case, the true values of $q$, $c$ and $\gamma_{0}$ need to be calculated via simulations since this distribution is not a member of the spherical family. It also follows that, differently from what observed for the multivariate Normal and $t$ cases, the values of these coefficients will be dependent on the chosen portfolio allocation. So they could change when moving from equal weighting to a different allocation scheme.
@@
For ease of reference, for each DGP, the true parameter values for $(a,b,q,\gamma_{0})$ are included in the True rows in Table A1.
The DCC-AL model is then fitted to each simulated dataset, using the estimation procedure as outlined in Section 3. The simulation results for the DCC-AL model are shown in Table A1 where, for ease of presentation, we have chosen not to report parameter estimates of the fitted first stage ES-CAViaR-IG models.
The rows labeled as True report, for each DGP, the values of the parameter true values used for simulation. The empirical averages of the 250 parameter estimates, for various return distributions and sample sizes, are shown in the Mean rows. The Root Mean Squared Error (RMSE) values between the parameter estimates and the true values are shown in the RMSE rows in Table A1.
@@
Finally, reminding that the main motivation for the DCC-AL model is the generation of accurate portfolio risk forecasts, the last two columns of Table A1 provide a very important benchmark for assessing the properties of the proposed estimation procedure. Comparing the $\alpha=2.5\%$ estimated and true risk forecasts, for both VaR and ES, it can be noted that these two series are on average very close even for the shortest sample size $T=2000$. The RMSE values are also remarkably low: with $T=2000$ they do not exceed 0.0811 for VaR and 0.1034 for ES, and their values monotonically decrease as $T$ increases across three return distributions. This last set of results confirms the ability of the proposed two-stage estimation procedure to accurately reproduce the portfolio risk dynamics.
Table 1: DCC-AL model parameter estimates and VaR and ES forecasting results with simulated datasets.
@@
$t$, $T=2000$ True 0.1200 0.7800 -0.5097 3.9717 -1.4353 -1.8159
Mean 0.1591 0.6719 -0.5844 4.1136 -1.4479 -1.8108
RMSE 0.1023 0.2419 0.1894 0.2967 0.0811 0.1034

### v4

# A semi-parametric dynamic conditional correlation framework for risk forecasting

@@ Abstract
We develop a novel multivariate semi-parametric framework for joint portfolio Value-at-Risk (VaR) and Expected Shortfall (ES) forecasting. Unlike existing univariate semi-parametric approaches, the proposed framework explicitly models the dependence structure among portfolio asset returns through a dynamic conditional correlation (DCC) parameterization. To estimate the model, a two-step procedure based on the minimization of a strictly consistent VaR and ES joint loss function is employed. This procedure allows to simultaneously estimate the DCC parameters and the portfolio risk factors. The performance of the proposed model in risk forecasting on various probability levels is evaluated by means of a forecasting study on the components of the Dow Jones index for an out-of-sample period from December 2016 to September 2021. The empirical results support effectiveness of the proposed framework compared to a variety of existing approaches.
Keywords: multivariate; semi-parametric; Value-at-Risk; Expected Shortfall; forecasting.
@@ 1 Introduction
Value-at-Risk (VaR) and Expected Shortfall (ES) play a central role in the risk management systems of banks and other financial institutions. For more than two decades, VaR has been the official risk measure adopted worldwide by financial intermediaries operating in the global financial system. Nevertheless, VaR has some important theoretical limits. First, VaR cannot measure the expected loss for extreme (violating) returns. In addition, it can be shown that VaR is not always a coherent risk measure, due to failure to match the subadditivity property. For these reasons, the Basel Committee on Banking Supervision proposed in May 2012 to replace VaR with the ES (Artzner 1997; Artzner et al. 1999). The prominent role of ES in the new international risk measurement practice is clearly illustrated by the official standards for determining minimum capital requirements for market risk published on the BIS institutional website (Basel Committee on Banking Supervision, 2019, see e.g. section MAR33) and effective from January 2022. Thus, in recent years ES has been increasingly employed for tail risk measurement. Analytically, ES is defined as the expectation of the return conditional on the VaR being exceeded and, unlike VaR, it is a coherent measure and “measures the riskiness of a position by considering both the size and the likelihood of losses above a certain confidence level” (Basel Committee on Banking Supervision, 2013).
Our paper focuses on the daily forecasting of VaR and ES on the lower/left tail. In line with the Basel III Accord and studies in the literature, our paper analyses the commonly used $\alpha=1\%$, 2.5% and 5% probability levels.
@@
We contribute to the literature on portfolio risk forecasting by proposing a novel class of semi-parametric multivariate GARCH (MGARCH) models that jointly generate portfolio VaR and ES forecasts, using multivariate information on the returns of portfolio constituents as input. Our main reference model is given by a semi-parametric version of the dynamic conditional correlations (DCC) model (Engle, 2002) via proposing a Semi-DCC framework. Its parameters are estimated by minimizing a strictly consistent joint VaR and ES loss function that fits into the FZ class of Fissler and Ziegel (2016), without requiring the formulation of restrictive assumptions on the conditional distribution of portfolio returns.
The proposed framework can be easily extended to other parameterizations from the multivariate GARCH literature such as the corrected DCC (cDCC) of Aielli (2013) or dynamic covariance models such as the BEKK model of Engle and Kroner (1995). Moreover, the Semi-DCC framework could be conveniently extended to more complex DCC parameterizations that allow for asymmetries and the inclusion of exogenous variables, such as the Asymmetric Generalized DCC (AG-DCC) family of models proposed by Cappiello et al. (2006).
@@
There are two different ways to look at our proposed framework. Namely, the Semi-DCC can be viewed either as a semi-parametric DCC model or as an ES-CAViaR for portfolio risk forecasting. It should be noted, however, that compared to each of these approaches, the Semi-DCC is characterized by some distinctive and innovative features.
@@
In fact, compared to the standard DCC, the structure of the Semi-DCC is made up of three components, of which the first two, relating to the specification of volatility and correlation dynamics, are common to the standard DCC model, while the third, relating to portfolio risk estimation, is specific to our model. We refer to our proposed model as a semi-parametric DCC model, since the risk model, while still using multivariate information on asset correlations, is fitted to the time series of univariate portfolio returns $r_{t}(\mathbf{w})$, for a given choice of portfolio weights $\mathbf{w}$. Furthermore, in the DCC model the dynamic conditional correlation parameters are usually estimated by Gaussian Quasi Maximum Likelihood (DCC-QML) through the maximization of a multivariate normal quasi-likelihood function. Then the portfolio VaR and ES could be ex-post obtained through the filtered historical simulation (HS) approach by calculating the sample quantiles and tail averages of the standardized returns (Francq and Zakoïan, 2015). In contrast, the Semi-DCC takes a more natural approach to VaR and ES estimation by fitting the risk model directly to the time series of portfolio returns rather than to the standardized returns. Specifically, the Semi-DCC involves the optimization of a univariate risk-targeted strictly consistent loss function for portfolio VaR and ES, e.g. the AL or FZ0 joint loss function, allowing portfolio VaR and ES to be estimated directly. An additional strength of this approach is that it also allows for the simultaneous estimation of dynamic correlation parameters and portfolio risk factors, whereas in standard DCC models these are estimated in two subsequent steps. The semi-nonparametric (SNP) DCC model proposed by Del Brio et al. (2011) and Níguez and Perote (2016) also replace the second-stage Gaussian QML with more flexible alternatives based on series expansions. Although this class of models is similar in spirit to ours in attempting to provide a more flexible alternative to the standard DCC, the Semi-DCC model differs from the SNP-DCC in two main ways. First, the Semi-DCC model is fitted to the series of univariate portfolio returns. Second, in the SNP-DCC the second stage estimation is based on an approximate likelihood that is not “risk-targeted” in the sense of being strictly consistent for the pair (VaR, ES).
@@
Compared to the standard ES-CAViaR framework, the Semi-DCC model extends it to a multivariate setting, where it allows for joint estimates of portfolio VaR and ES that explicitly take into account the cross-sectional correlation structure of portfolio returns. From a risk forecasting perspective, the Semi-DCC model can then be viewed as an ES-CAViaR that takes multivariate input information on the asset variance and covariance matrix, which is then used to obtain univariate VaR and ES measures for portfolio returns. However, unlike univariate semi-parametric approaches to risk forecasting, the underlying multivariate GARCH structure allows the Semi-DCC model to be used for a wider range of applications, including portfolio allocation and hedging.
@@
The empirical performance of the Semi-DCC using AL and FZ0 joint losses model is assessed by means of an application to a panel of 28 assets included in the Dow Jones index. Our findings can be summarized as follows. When forecasting risk for an equally weighted portfolio, Semi-DCC models are competitive with state-of-the-art univariate semi-parametric approaches and perform better than conventional DCC models. Further, we assess the performance of the proposed model in constructing global minimum ES portfolios. Under this respect, our results show that the Semi-DCC model clearly outperforms a benchmark equally weighted allocation strategy. We also find evidence that in a turbulent period centred on the outbreak of the COVID-19 crisis in 2020, the Semi-DCC model leads to less risky portfolios compared to existing DCC specifications.
The paper is structured as follows. Section 2, after providing a brief review of the existing semi-parametric univariate risk forecasting approaches, presents the proposed class of semi-parametric DCC models, while their estimation procedure is illustrated in Section 3. The finite sample properties of the estimators and the risk forecasting performance of Semi-DCC are investigated in Section 4 via an extensive Monte Carlo simulation study. Section 5 discusses the results of an empirical application to risk forecasting for the constituents of the Dow Jones index, where the portfolio optimization results are included in Appendix A to limit the focus of the empirical study. Section 6 concludes the paper.
@@ 2.1 Description of the environment
In the remainder of Sections 2 and 3, considering that we are studying risk forecasting at a given probability level (1%, 2.5% or 5%) for a pre-defined portfolio composition $(\mathbf{w})$, the following notational conventions will be adopted to simplify the presentation: $Q_{t,\alpha}(\mathbf{w})\equiv Q_{t}$, $\text{ES}_{t,\alpha}(\mathbf{w})\equiv\text{ES}_{t}$, $q_{\alpha}(\mathbf{w})\equiv q$, $c_{\alpha}(\mathbf{w})\equiv c$ and $r_{t}(\mathbf{w})\equiv r_{t}$.
@@ 2.2 Univariate semi-parametric approaches to portfolio risk forecasting
This allows a quasi-likelihood function to be built and maximised, given model expressions for $(Q_{t},\,\text{ES}_{t})$. Taylor (2019) supports the validity of this estimation procedure noting that the negative logarithm of the resulting likelihood function is strictly consistent for $(Q_{t},\,\text{ES}_{t})$ considered jointly, that is it fits into the FZ class of jointly consistent scoring functions for VaR and ES developed by Fissler and Ziegel (2016).
Along the same lines, Patton et al. (2019) investigate a class of semi-parametric models, including some observation driven models, whose parameters can be estimated minimizing a 0-degree homogeneous loss function (FZ0) still included in the same FZ class, assuming that both VaR and ES are both strictly negative. The FZ0 loss used in Patton et al. (2019) can be shown as:
$FZ0=-\frac{1}{\alpha\text{ES}_{t}}I(r_{t}\leq Q_{t})\,(Q_{t}-r_{t})+\frac{Q_{t}}{\text{ES}_{t}}+\log(-\text{ES}_{t})-1$ (5)
As mentioned above, all these papers focus on univariate semi-parametric modeling approaches that, when applied to portfolio returns, do not explicitly assess the impact of cross-sectional correlations among assets. Although this issue has been extensively analyzed in the literature on parametric MGARCH models (Bauwens et al., 2006), there is much less existing work in addressing the issue in a multivariate semi-parametric risk-targeted framework. This motivates the proposal of a novel semi-parametric DCC modeling strategy.
2.3 A semi-parametric DCC framework
The analytical expressions for portfolio VaR and ES provided in Equation (3) make evident that the VaR and ES dynamics are driven by those of the conditional variance and covariance matrix $H_{t}$. In this section we describe a semi-parametric framework for assessing the impact that conditional correlations among portfolio components and their individual volatilities have on portfolio risk forecasts.
The dynamics of $H_{t}$ can be modelled using a wide range of specifications from the MGARCH literature. Without any loss of generality, due to its flexibility and widespread diffusion among practitioners, our proposed modeling approach builds on the DCC model of Engle (2002) and shares the same volatility and correlation dynamics. However, compared to the standard DCC, our framework incorporates an additional step related to the specification of portfolio VaR and ES.
In particular, the risk model is fitted to the time series of univariate portfolio returns $r_{t}(\mathbf{w})$, for a given set of portfolio weights $\mathbf{w}$. The semi-parametric nature of the model derives from the fact that, as it will be later discussed in Section 3, its coefficients are fitted minimizing a jointly consistent loss for $Q_{t}$ and $\text{ES}_{t}$, e.g., AL (Equation (4)) or FZ0 (Equation (5)), without assuming the parametric distribution of returns.
@@ 2.3 A semi-parametric DCC framework
Given the assumptions on $a$ and $b$, no identifiability issue arises and only four coefficients need to be estimated in the second stage of the Semi-DCC model: $q$, $a$, $b$ and $\gamma_{0}$. The parameter estimate $q$ (VaR factor) allows us to directly estimate the quantile of the error distribution in the model for portfolio returns. Since we have defined $c^{2}/q^{2}=1+\text{exp}(\gamma_{0})$, after estimating parameters $q$ and $\gamma_{0}$, the value of the ES factor $c$ can be immediately recovered. The VaR and ES joint loss function from the FZ class can be incorporated into the proposed Semi-DCC for model estimation. In this paper, we focus on testing AL and FZ0 losses, with the corresponding namings of Semi-DCC-AL and Semi-DCC-FZ0 used.
@@ 3 Estimation
The estimation of the Semi-DCC model can be performed by means of a two step procedure in the spirit of Engle (2002). For demonstration purpose, the Semi-DCC using the AL loss is presented in this section, while other VaR and ES joint loss functions such as FZ0 can be directly incorporated, by replacing AL with FZ0 as the optimization target.
@@
In step 1, we use the Matlab 2023a “fmincon” optimization routine to estimate the ES-CAViaR-IG models. In step 2, the Matlab “MultiStart” facility is further employed for “fmincon” to improve the robustness of the estimation results. We use 5 separate sets of starting points, generating 5 local solutions, then the optimum set among these is finally chosen as the parameter estimates. The sensitivity of strictly consistent loss functions, such as AL and FZ0, to initial conditions is a well-known problem that typically complicates the estimation of model parameters in semi-parametric risk models (Engle and Manganelli, 2004). Nevertheless, estimation strategies based on the use of multiple initial values, such as the one implemented in this paper, have been found in the literature to be effective in overcoming this sensitivity (see Taylor, 2019, among others).
@@ 4 Simulation study
A simulation study has been conducted in order to assess the statistical properties of the two-step VaR and ES joint loss based estimation procedure discussed in Section 3. We focus on presenting the simulation results of Semi-DCC-AL when the AL based loss is used, since we find that the performances of Semi-DCC-AL and Semi-DCC-FZ0 are very close. In the following empirical section, risk forecasting results of both Semi-DCC-AL and Semi-DCC-FZ0 will be presented.
Our simulation analysis has two main objectives. First, the study evaluates the bias and efficiency of the estimators of the Semi-DCC-AL parameters ($a,b,q,\gamma_{0}$). Second, we assess the the risk forecasting performances of the estimated Semi-DCC-AL model and its comparison to parametric DCC models in an equally weighted portfolio, via comparing the one-step-ahead level portfolio VaR and ES forecast accuracy to the “true” simulated values.
4.1 Semi-DCC-AL performance
@@
On the $\alpha=2.5\%$ probability level [^4] We have also tested the $1\%$ and $5\%$ probability levels, while the observations on the Semi-DCC-AL parameter estimates and risk forecasts are in general consistent with the ones from $2.5\%$, thus only the 2.5% results are presented in this section. In the following section, we compare the risk forecasting performance of the standard parametric DCC models with the Semi-DCC-AL on the $1\%$, $2.5\%$ and $5\%$ levels in detail., the Semi-DCC-AL model is then fitted to the time series of equally weighted $(\mathbf{w}=1/n)$ portfolio returns generated from the chosen DGP, for three different sample sizes $T\in\{2000,3000,5000\}$. Overall, matching 3 different distributional assumptions (including both spherical and non-spherical distributions) and sample sizes, 9 simulation settings have been obtained and, under each of these, 250 return series have been generated.
Conditional on the chosen error distribution and on the values of the correlation parameters $a$ and $b$, the true values of the other two parameters $q$ and $\gamma_{0}$ in the Semi-DCC-AL model can be then calculated as follows.
@@
Therefore, taking the multivariate Normal case on the $\alpha=2.5\%$ probability as an example, in the Semi-DCC-AL the true value of $q^{2}$ is $\Phi^{-1}(\alpha)^{2}=(-1.96)^{2}=3.8415$. The true value for $c^{2}$ is $\left(\frac{\phi(\Phi^{-1}(\alpha))}{\alpha}\right)^{2}=(-2.3378)^{2}$, where $\Phi$ is the standard Normal CDF and $\phi$ is standard Normal Probability Density Function (PDF). Further, we have $(1+\text{exp}(\gamma_{0}))=\frac{c^{2}}{q^{2}}=\frac{(-2.3378)^{2}}{(-1.9600)^{2}}=1.4227$, thus $\gamma_{0}=\text{log}(1.4227-1)=-0.8610$. These true values of $a$, $b$, $\gamma_{0}$ and $q^{2}$ for the multivariate Normal distribution are shown in the True rows in Table 1. When $\mathbf{z}_{t}\overset{iid}{\sim}t_{n}(\mathbf{0},I_{n};\nu=10)$, the true values of $q$, $c$ and $\gamma_{0}$ are obtained through a similar approach, replacing the analytical formulas for $q$ and $c$ with the equivalent expressions for an univariate $t$ distribution.
In the $nst$ case, the true values of $q$, $c$ and $\gamma_{0}$ need to be calculated via simulations since this distribution is not a member of the spherical family. It also follows that, differently from what observed for the multivariate Normal and $t$ cases, the values of these coefficients will be dependent on the chosen portfolio allocation, thus they could change when moving from equal weighting to a different allocation scheme.
@@
For ease of reference, for each DGP, the true parameter values for $(a,b,q,\gamma_{0})$ are included in the True rows in Table 1.
The Semi-DCC-AL model is then fitted to each simulated dataset, using the estimation procedure as outlined in Section 3. The simulation results for the Semi-DCC-AL model are shown in Table 1 where, for ease of presentation, we have chosen not to report parameter estimates of the fitted first stage ES-CAViaR-IG models.
The rows labeled as True report, for each DGP, the values of the parameter true values used for simulation. The empirical averages of the 250 parameter estimates, for various return distributions and sample sizes, are shown in the Mean rows. The Root Mean Squared Error (RMSE) values between the parameter estimates and the true values are shown in the RMSE rows in Table 1.
@@
Finally, reminding that the main motivation for the Semi-DCC-AL model is the generation of accurate portfolio risk forecasts, the last two columns of Table 1 provide a very important benchmark for assessing the properties of the proposed estimation procedure. Comparing the $\alpha=2.5\%$ estimated and true risk forecasts, for both VaR and ES, it can be noted that these two series are on average very close even for the shortest sample size $T=2000$. The RMSE values are also remarkably low: with $T=2000$ they do not exceed 0.0811 for VaR and 0.1034 for ES, and their values monotonically decrease as $T$ increases across three return distributions. This last set of results confirms the ability of the proposed two-stage estimation procedure to accurately reproduce the portfolio risk dynamics.
Table 1: Semi-DCC-AL model parameter estimates and VaR and ES forecasting results with simulated datasets.
@@
$t$, $T=2000$ True 0.1200 0.7800 -0.5097 3.9717 -1.4361 -1.8169
Mean 0.1591 0.6766 -0.5839 4.1157 -1.4499 -1.8135
RMSE 0.1033 0.2437 0.1894 0.2954 0.0818 0.1045